OpenStudy (anonymous):

Does anyone know how to do this? Mr.Jenkins can dig a ditch in one hour less time than his son. Working together, they can do the job in 2/3 hour. How long does it take Mr.Jenkins to do the job alone? The answer is one huor

OpenStudy (anonymous):

digging time for the son S, then S - 1 is the digging time for the dad (1/S + 1/(S-1))(2/3) = 1 S = 2hrs so F = 2-1 = 1hr

OpenStudy (whpalmer4):

There's also another solution to the equation, s = 1/2 hour, but that clearly doesn't work if dad is going to take an hour less!